# ModelNet40-E: An Uncertainty-Aware Benchmark for Point Cloud Classification

### v1

@@ Abstract
We introduce ModelNet40-E, a new benchmark designed to assess the robustness and calibration of point cloud classification models under synthetic LiDAR-like noise. Unlike existing benchmarks, ModelNet40-E provides both noise-corrupted point clouds and point-wise uncertainty annotations via Gaussian noise parameters ($\sigma$, $\mu$), enabling fine-grained evaluation of uncertainty modeling. We evaluate three popular models—PointNet, DGCNN, and Point Transformer v3—across multiple noise levels using classification accuracy, calibration metrics, and uncertainty-awareness. While all models degrade under increasing noise, Point Transformer v3 demonstrates superior calibration, with predicted uncertainties more closely aligned with the underlying measurement uncertainty.
Introduction
@@
In recent years, 3D point cloud classification has gained increasing importance across a wide range of applications, including autonomous driving (Lang et al. 2018; Yan, Mao, and Li 2018; Sun et al. 2019), robotics (Mousavian, Eppner, and Fox 2019), augmented reality (Izadi et al. 2011; Newcombe, Fox, and Seitz 2015; Jiang et al. 2025; Newcombe et al. 2011; Yu et al. 2020), and remote sensing (Wang et al. 2025, 2024; Pan et al. 2025). In response, numerous models have been proposed, ranging from early architectures like PointNet (Qi et al. 2017a) and PointNet++ (Qi et al. 2017b) to more advanced designs such as DGCNN (Wang et al. 2019), Point2Vec (Abou Zeid et al. 2023), and Point Transformer (Zhao et al. 2021; Wu et al. 2022, 2024).
@@
Constrained by the limited availability of real-world annotated point cloud data, most of these models are trained and evaluated on synthetic datasets, such as ModelNet40 (Wu et al. 2015) and Shapenet (Chang et al. 2015). These datasets are composed of clean, idealized point clouds that lack the noise, occlusions, and distortions typically present in real sensor data. As a result, models that perform well in these benchmarks often fail to generalize to real-world environments (Wu et al. 2018; Zhao, Zhao, and Lee 2022; Ren, Yang, and Velipasalar 2025).
@@
To ensure fair comparisons and improve the applicability of the models in real-world scenarios, robustness benchmarks like ModelNet-C (Sun et al. 2022) have been proposed. These benchmarks introduce synthetic corruptions such as rotations, jittering, scaling, and point dropout to test the resilience of models to input perturbations. Despite offering valuable insights, they face key limitations: (1) they apply generic perturbations that do not reflect realistic LiDAR noise, and (2) they lack ground-truth uncertainty annotations. In safety-critical applications, such as autonomous driving and robotic manipulation, it is not enough for a model to simply output a class label. Equally important is assessing how confident the model is in its predictions. Overconfident yet incorrect predictions can lead to catastrophic failures, making calibration and uncertainty awareness essential attributes for building reliable models.
@@
We introduce ModelNet40-E, a novel benchmark that builds upon the original ModelNet40 dataset by corrupting it with LiDAR-like noise at three severity levels: Light, Moderate, and Heavy. Unlike random jitter or dropout, LiDAR noise is range-dependent, anisotropic, and often structured, influenced by surface reflectivity, incidence angles, and environmental conditions. Capturing these complexities is critical for creating realistic benchmarks that truly test model robustness and uncertainty estimation. Our noise simulation is inspired by empirical studies of LiDAR measurement errors (Gschwandtner et al. 2011), incorporating range-dependent Gaussian noise, incidence angle effects, systematic bias, and random outliers. Crucially, for every point in the dataset, we also provide the corresponding per-point noise statistics: standard deviation $\sigma$ and expected bias $\mu$. This allows for rigorous evaluation of both predictive performance and uncertainty estimation.
@@
  • We introduce ModelNet40-E, a realistic, uncertainty-aware benchmark for point cloud classification.
  • We evaluate three popular point cloud models—PointNet, DGCNN, and PTv3—across several metrics for different noise levels.
  • We show that while all models degrade under noise, PTv3 demonstrates superior robustness and calibration, particularly in aligning predicted confidence with ground-truth uncertainty.
Related Work
Point Cloud Classification Benchmarks
[…]
@@
Robustness and Adversarial Attacks on Point Clouds
Robustness in point cloud classification refers to a model’s ability to maintain performance under various perturbations. These include sensor noise, occlusions, point sparsity, and misalignment, which are common challenges present in real-world data. To evaluate resilience under such conditions, recent benchmarks like ModelNet40-C (Sun et al. 2022) apply a wide range of synthetic corruptions (e.g., rotation, scaling, dropout, jitter) and measure performance across different perturbation levels.
Beyond natural corruptions, there has been growing interest in studying adversarial robustness, where small, carefully crafted perturbations are designed to fool classifiers. Several works have extended adversarial attack strategies from the 2D image domain to 3D point clouds. For instance, Xiang, Qi, and Li (2018) proposes adversarial perturbation and point generation strategies to mislead PointNet into high-confidence misclassifications. In addition, Zheng et al. (2018) introduces a point perturbation attack that uses gradient information to identify and remove critical points, degrading classification accuracy. Other approaches, like Liu, Yu, and Su (2019) and Tsai et al. (2020), have explored transformations and translations to mislead classifiers while preserving overall object shape.
To counter these attacks, a variety of defense strategies have been developed, including adversarial training, which improves model robustness by exposing it to adversarial examples during training (Sun et al. 2020), as well as input reconstruction methods that denoise or repair corrupted point clouds before classification (Zhou et al. 2018). However, many of these defenses come at the cost of reduced clean accuracy, and few have been evaluated under realistic sensor noise conditions, limiting their practical effectiveness.
@@
Uncertainty Estimation in Deep Learning
Uncertainty estimation in deep learning refers to a model’s ability to not only make predictions but also to quantify how confident it is in those predictions. In a well-calibrated model, confidence scores should decrease appropriately as input perturbations are introduced. If confidence scores remain high even in the presence of significant corruptions, this may lead to overconfident misclassifications. It is well established that neural networks are often miscalibrated and tend to be overconfident, particularly as model capacity or training time increases (Guo et al. 2017).
@@
Several techniques exist to estimate these uncertainties. For example, Gal and Ghahramani (2016) proposes the Monte Carlo (MC) dropout method, which performs multiple stochastic forward passes with dropout enabled at test time to approximate predictive uncertainty. In addition, deep ensembles (Lakshminarayanan, Pritzel, and Blundell 2017) train multiple models with different initial conditions and estimate uncertainty by measuring the variance between their outputs.
ModelNet40-E
Benchmark Overview
ModelNet40-E is an uncertainty-aware benchmark for point cloud classification. It is built upon the original ModelNet40 dataset (Wu et al. 2015), introducing several levels of LiDAR-like noise based on realistic sensor noise. By providing per-point noise statistics ($\sigma$ and $\mu$), the main goal of ModelNet40-E is to serve as a common benchmark for evaluating the uncertainty estimation capabilities of point cloud classification models.
@@
LiDAR Noise Simulation
To approximate realistic sensor noise, we simulate a LiDAR acquisition process inspired by empirical studies of LiDAR measurement errors (Gschwandtner et al. 2011). For each clean point cloud, we compute per-point range and incidence angle relative to the sensor. Measurement noise is sampled from a Gaussian distribution whose standard deviation increases with distance and varies systematically with surface orientation. Specifically, the per-point noise $\epsilon$ is sampled as
@@
Bias ($k$). In addition to the range-dependent noise arising from increasing measurement uncertainty with distance, there is another contribution from systematic sensor calibration offset. This effect also depends on range and is defined as
@@
where $k$ is a deterministic bias term added to the measured range.
Outlier probability. Outliers are common in LiDAR data, often caused by spurious reflections, multipath returns, or transient measurement artifacts. We define an outlier probability parameter specifying the proportion of points that are replaced with random points—sampled uniformly within a bounding box around the object—during ModelNet40-E generation.
@@
Table 1: Configuration parameters for the light, moderate, and heavy noise levels in ModelNet40-E.
Tier Range Noise $(a,b)$ Angle Factor $c$ Bias $k$ Outlier probability
Light (0.003, 0.001) 1.5 0.005 0.01
Moderate (0.005, 0.002) 2.0 0.010 0.02
Heavy (0.010, 0.003) 3.0 0.015 0.05
Figure 1: Examples of ModelNet40 point clouds under different noise levels: none, light, moderate, and heavy. ”none” refers to the original, unperturbed point cloud.
Experiments
Experimental Setup
We evaluate three representative architectures—PointNet, DGCNN, and PTv3—on ModelNet40-E. Each model is trained independently on each noise level using the training set and evaluated on the corresponding test set. We report classification accuracy, calibration error, and uncertainty correlation.
Models
(1) PointNet, introduced by Qi et al. (2017a), was the first end-to-end architecture to directly process raw point clouds without requiring meshing or voxelization. It follows the next steps: (1) per-point feature extraction via shared multilayer perceptrons (MLPs), and (2) a max-pooling operation that aggregates the per-point features into a single global feature vector. This design allows the model to process unordered point sets and capture global shape information.
For classification tasks, the global feature vector is passed through a series of three fully connected (FC) layers that form the classification head. Each FC layer is followed by a ReLU activation and dropout, except for the last layer.
(2) Dynamic Graph CNN (DGCNN; Wang et al. 2019) is a point cloud classification and segmentation model that captures local geometric features through a key operation called EdgeConv. It starts by constructing a graph for each point based on its k-nearest neighbors (kNN) in feature space. For each edge, a local feature is computed by applying a multilayer perceptron (MLP) to the concatenation of the point’s feature and the relative offset between the point and its neighbors.
A distinctive feature of DGCNN is its dynamic graph construction: after each layer, the neighborhood graph is updated according to the new point features rather than their coordinates.
Finally, the model aggregates features across layers and uses a global pooling followed by fully connected (FC) layers to perform classification or segmentation.
(3) Point Transformer v3 (PTv3; Wu et al. 2024) is a transformer-based architecture designed for point cloud understanding. The original Point Transformer (PT; Zhao et al. 2021) first introduced attention mechanisms to capture local geometric structures in point clouds, and Point Transformer v2 (PTv2; Wu et al. 2022) further improved scalability and efficiency. PTv3 builds on these ideas, operating directly on unordered point clouds with multiple stages of attention blocks that capture both local and global dependencies. Using learned attention weights, each block aggregates contextual information by attending to nearby points.
In our implementation, we use PTv3 with only the $(x,y,z)$ coordinates as input and FlashAttention disabled. The network consists of five stages of transformer blocks followed by a max-pooling operation. A simple classification head, composed of two fully connected layers, outputs the predicted class probabilities for each sample.
Noise Configurations
We employ four different noise configurations: None, Light, Moderate, and Heavy. ”None” refers to the original, unperturbed ModelNet40 dataset. The parameters used to generate the noise are summarized in Table 1. For this benchmark, the virtual sensor position is fixed at (0,2,0)(0,-2,0)( 0 , - 2 , 0 ), which defines the range and incidence angles used in the noise simulation.
Figure 1 illustrates examples of point clouds for each noise level.
Training Details
Models are trained independently on each noise configuration (None, Light, Moderate, Heavy) for 100 epochs using cross-entropy loss as the objective. Training is performed with the Adam optimizer (Kingma 2014), using an initial learning rate of $1\times 10^{-3}$, which is decayed by a factor of 0.7 every 20 epochs. PointNet and DGCNN models are trained with a batch size of 32, while PTv3 uses a reduced batch size of 8 due to its higher memory requirement. The best models are selected based on the highest classification accuracy achieved on the test set. All experiments were conducted on a single NVIDIA RTX 4090 GPU.
Evaluation Metrics
@@
(1) Classification Accuracy. This metric evaluates the predictive performance of classification models. It is defined as the proportion of correctly predicted test samples. In our experiments, we report accuracy in the test set for each noise level independently. It is expected that accuracy decreases as the noise level increases, and a lower decrement signals better robustness to noise.
[…]
@@
(3) Correlation Between Measurement Uncertainty and Predicted Uncertainty. We evaluate whether models can accurately recognize the reliability of their predictions. This is an essential property for avoiding overconfident or underconfident outputs. To assess this, we compute the Pearson correlation between the ground truth measurement uncertainty of each sample and the model’s predicted uncertainty.
For each sample, we define the predicted uncertainty as one minus the predicted confidence, $1-p_{\text{max}}$, where $p_{\text{max}}$ is the maximum softmax probability assigned to any class. Ground truth measurement uncertainty is quantified by the mean standard deviation $\sigma$ of the simulated noise applied to that sample.
A higher Pearson correlation indicates that the model’s confidence decreases as the input noise increases, demonstrating better awareness of input uncertainty.
Main Results
[…]
Classification Accuracy
Table 2 presents the classification accuracy on the test set for all models across different noise levels. As expected, performance degrades as the noise level increases. However, this decline is notably more pronounced in PointNet compared to DGCNN and PTv3, indicating that PointNet is less robust to corruption.
Table 2: Classification Accuracy (%) for all models across different noise levels.

### v2

@@ Abstract
Most existing benchmarks for point cloud classification focus solely on accuracy, overlooking critical aspects such as calibration and uncertainty awareness that are essential for safety-critical applications. We introduce ModelNet40-E, a benchmark that complements prior robustness efforts by providing noise-corrupted point clouds together with per-point uncertainty annotations via Gaussian parameters ($\sigma$, $\mu$). Unlike benchmarks based on random corruptions, ModelNet40-E introduces physically motivated LiDAR-like noise at multiple levels, reflecting real-world sensing conditions. Using this benchmark, we evaluate a range of representative point cloud architectures across varying noise levels. Our results show that accuracy alone can be misleading: some models with lower accuracy exhibit better calibration and uncertainty awareness, underscoring the need to evaluate all three metrics together.
1 Introduction
@@
In recent years, 3D point cloud classification has gained increasing importance across a wide range of applications, including autonomous driving (Lang et al., 2018; Yan et al., 2018; Sun et al., 2019), robotics (Mousavian et al., 2019), augmented reality (Izadi et al., 2011; Newcombe et al., 2015; Jiang et al., 2025; Newcombe et al., 2011; Yu et al., 2020), and remote sensing (Wang et al., 2025; 2024; Pan et al., 2025). In response, numerous models have been proposed, ranging from early architectures like PointNet (Qi et al., 2017a) and PointNet++ (Qi et al., 2017b) to more advanced designs such as DGCNN (Wang et al., 2019), Point2Vec (Abou Zeid et al., 2023), and Point Transformer (Zhao et al., 2021; Wu et al., 2022; 2024).
@@
Constrained by the limited availability of real-world annotated point cloud data, most of these models are trained and evaluated on synthetic datasets, such as ModelNet40 (Wu et al., 2015) and Shapenet (Chang et al., 2015). These datasets are composed of clean, idealized point clouds that lack the noise, occlusions, and distortions typically present in real sensor data. As a result, models that perform well in these benchmarks often fail to generalize to real-world environments (Wu et al., 2018; Zhao et al., 2022; Ren et al., 2025).
@@
To ensure fair comparisons and improve the applicability of the models in real-world scenarios, robustness benchmarks like ModelNet-C (Sun et al., 2022) have been proposed. These benchmarks introduce synthetic corruptions such as rotations, jittering, scaling, and point dropout to test the resilience of models to input perturbations. Despite offering valuable insights, they face key limitations: (1) they apply generic perturbations that do not reflect realistic LiDAR noise, and (2) they lack ground-truth uncertainty annotations. In safety-critical applications, such as autonomous driving and robotic manipulation, it is not enough for a model to simply output a class label. Equally important is assessing how confident the model is in its predictions. Overconfident yet incorrect predictions can lead to catastrophic failures, making calibration and uncertainty awareness essential attributes for building reliable models.
@@
We introduce ModelNet40-E, a novel benchmark that builds upon the original ModelNet40 dataset by corrupting it with LiDAR-like noise at three severity levels: Light, Moderate, and Heavy. Unlike random jitter or dropout, LiDAR noise is range-dependent, anisotropic, and often structured, influenced by surface reflectivity, incidence angles, and environmental conditions. Capturing these complexities is critical for creating realistic benchmarks that truly test model robustness and uncertainty estimation. Our noise simulation is inspired by empirical studies of LiDAR measurement errors (Gschwandtner et al., 2011), incorporating range-dependent Gaussian noise, incidence angle effects, systematic bias, and random outliers. Crucially, for every point in the dataset, we also provide the corresponding per-point noise statistics: standard deviation $\sigma$ and expected bias $\mu$. This allows for rigorous evaluation of both predictive performance and uncertainty estimation.
@@
  • We introduce ModelNet40-E, a benchmark for point cloud classification that extends the standard ModelNet40 dataset with realistic LiDAR-inspired noise, enabling evaluation not only of classification accuracy but also of calibration, and uncertainty awareness.
  • We propose a LiDAR-like noise generation procedure that injects range-dependent, angle-dependent, and outlier noise, providing controlled yet realistic perturbations for benchmarking robustness.
  • Using ModelNet40-E, we conduct a systematic evaluation of several representative architectures, highlighting that while performance degrades across noise levels, models differ significantly in robustness, calibration, and error detection ability.
2 Related Work
2.1 Point Cloud Classification Benchmarks
3D point cloud classification models are commonly evaluated on synthetic datasets such as ModelNet40 (Wu et al., 2015) and Shapenet (Chang et al., 2015). Our work focuses on ModelNet40, a widely used benchmark consisting of 12,311 CAD-generated point clouds across 40 categories. The dataset is split into training ($80\%$) and testing ($20\%$) sets, containing 9,840 and 2,468 samples, respectively. It contains clear and uniformly sampled point clouds, making it easily processed and suitable as a standard benchmark.
@@ 2.1 Point Cloud Classification Benchmarks
2.2 Robustness and Adversarial Attacks on Point Clouds
Robustness in point cloud classification refers to a model’s ability to maintain performance under various perturbations. These include sensor noise, occlusions, point sparsity, and misalignment, which are common challenges present in real-world data. To evaluate resilience under such conditions, recent benchmarks like ModelNet40-C (Sun et al., 2022) apply a wide range of synthetic corruptions (e.g., rotation, scaling, dropout, jitter) and measure performance across different perturbation levels.
Beyond natural corruptions, there has been growing interest in studying adversarial robustness, where small, carefully crafted perturbations are designed to fool classifiers. Several works have extended adversarial attack strategies from the 2D image domain to 3D point clouds. For instance, Xiang et al. (2018) proposes adversarial perturbation and point generation strategies to mislead PointNet into high-confidence misclassifications. In addition, Zheng et al. (2018) introduces a point perturbation attack that uses gradient information to identify and remove critical points, degrading classification accuracy. Other approaches, like Liu et al. (2019) and Tsai et al. (2020), have explored transformations and translations to mislead classifiers while preserving overall object shape.
To counter these attacks, a variety of defense strategies have been developed, including adversarial training, which improves model robustness by exposing it to adversarial examples during training (Sun et al., 2020), as well as input reconstruction methods that denoise or repair corrupted point clouds before classification (Zhou et al., 2018). However, many of these defenses come at the cost of reduced clean accuracy, and few have been evaluated under realistic sensor noise conditions, limiting their practical effectiveness.
@@ 2.2 Robustness and Adversarial Attacks on Point Clouds
2.3 Uncertainty Estimation in Deep Learning
Uncertainty estimation in deep learning refers to a model’s ability to not only make predictions but also to quantify how confident it is in those predictions. In a well-calibrated model, confidence scores should decrease appropriately as input perturbations are introduced. If confidence scores remain high even in the presence of significant corruptions, this may lead to overconfident misclassifications. It is well established that neural networks are often miscalibrated and tend to be overconfident, particularly as model capacity or training time increases (Guo et al., 2017).
@@ 2.3 Uncertainty Estimation in Deep Learning
Several techniques exist to estimate these uncertainties. For example, Gal & Ghahramani (2016) proposes the Monte Carlo (MC) dropout method, which performs multiple stochastic forward passes with dropout enabled at test time to approximate predictive uncertainty. In addition, deep ensembles (Lakshminarayanan et al., 2017) train multiple models with different initial conditions and estimate uncertainty by measuring the variance between their outputs.
3 ModelNet40-E
3.1 Benchmark Overview
ModelNet40-E is an uncertainty-aware benchmark for point cloud classification. It is built upon the original ModelNet40 dataset (Wu et al., 2015), introducing several levels of LiDAR-like noise based on realistic sensor noise. By providing per-point noise statistics ($\sigma$ and $\mu$), the main goal of ModelNet40-E is to serve as a common benchmark for evaluating the uncertainty estimation capabilities of point cloud classification models.
@@ 3.1 Benchmark Overview
3.2 LiDAR Noise Simulation
To approximate realistic sensor noise, we simulate a LiDAR acquisition process inspired by empirical studies of LiDAR measurement errors (Gschwandtner et al., 2011). For each clean point cloud, we compute per-point range and incidence angle relative to the sensor. Measurement noise is sampled from a Gaussian distribution whose standard deviation increases with distance and varies systematically with surface orientation. Specifically, the per-point noise $\epsilon$ is sampled as
@@ 3.2 LiDAR Noise Simulation
Bias ($k$). In addition to range- and angle-dependent measurement noise, we include a systematic bias term related to sensor calibration. This bias is modeled as
@@
where $\theta$ is the incidence angle and $k$ controls the magnitude of the offset.
Outlier probability. Outliers are common in LiDAR data, often caused by spurious reflections, multipath returns, or transient measurement artifacts. We define an outlier probability parameter specifying the proportion of points that are replaced with random points sampled uniformly within the normalized bounding box $0.5,0.5]\times[-0.5,0.5]\times[-0.5,0.5]$ around the object.
@@
3.3 Noise Configurations
We consider four noise configurations: None, Light, Moderate, and Heavy. The None setting corresponds to the original, unperturbed ModelNet40 dataset. For the corrupted versions, the parameters are sampled from the ranges summarized in Table 1. During noise simulation, the virtual sensor position is randomly sampled on a sphere of radius 2.0, with azimuth uniformly drawn from $0,2\pi]$ and elevation from $\pi/4,\pi/4]$.
Table 1: Configuration parameters for the light, moderate, and heavy noise levels in ModelNet40-E. For each point cloud, the parameters $(a,b)$, $c$, $k$, and the outlier probability are sampled uniformly from the ranges shown in the table.
@@ 3.3 Noise Configurations
Light ([2.0, 4.0], [0.5, 1.5]) [1.0, 2.0] [2.5, 7.5] [0.5, 1.5]
Moderate ([3.0, 7.0], [1.0, 3.0]) [1.5, 2.5] [5.0, 15.0] [1.0, 3.0]
Heavy ([5.0, 15.0], [2.0, 4.0]) [2.0, 4.0] [10.0, 25.0] [4.0, 8.0]
4 Experiments
4.1 Experimental Setup
We evaluate our benchmark on several representative architectures for point clouds: PointNet (Qi et al., 2017a), PointNet++ (Qi et al., 2017b), DGCNN (Wang et al., 2019), SimpleView (Goyal, 2020), CurveNet (Xiang et al., 2021), PointMLP (Ma et al., 2022), and Point Transformer v3 (PTv3; Wu et al. 2024). Each model is trained independently on the original ModelNet40 training dataset (”None” noise level in our paper), and evaluated on each noise level (”None”, ”Light”, ”Moderate”, and ”Heavy”) test set. We report classification accuracy, Expected Calibration Error (ECE), AUROC (Error Detection), and Pearson correlation between ground-truth measured and predicted uncertainty.
All models are trained for 100 epochs using cross-entropy loss and the Adam optimizer (Kingma, 2014), with an initial learning rate of $1\times 10^{-3}$ decayed by a factor of 0.7 every 20 epochs. The batch size is 32 for all models except PTv3, which uses a reduced batch size of 8 due to its higher memory requirements. To ensure fair comparison across accuracy, calibration, and uncertainty awareness, we consistently evaluate the checkpoint from the final epoch (epoch 100) rather than selecting models based on accuracy alone. All experiments were conducted on a single NVIDIA RTX 4090 GPU.
4.2 Evaluation Metrics
Figure 1: Calibration curves for all models across different noise levels. The dashed diagonal line represents perfect calibration. The gray shaded area indicates confidence values under $40\%$, which are underrepresented and less reliable.
4.2.1 Classification Accuracy
Accuracy evaluates the predictive performance of classification models. It is defined as the proportion of correctly predicted test samples. In our experiments, we report accuracy in the test set for each noise level independently. It is expected that accuracy decreases as the noise level increases, and a lower decrement signals better robustness to noise.
4.2.2 Expected Calibration Error (ECE)
ECE (Hackel et al., 2017; Naeini et al., 2015; Nixon et al., 2019) quantifies the discrepancy between a model’s predicted confidence and its actual accuracy. In a perfectly calibrated model, the predicted confidence matches the observed accuracy. For instance, predictions made with $70\%$ confidence should be correct approximately $70\%$ of the time.
@@ 4.2.2 Expected Calibration Error (ECE)
4.2.3 Uncertainty Awareness
We evaluate uncertainty awareness using two complementary measures.
(1) AUROC. AUROC is a standard metric that evaluates the model’s ability to detect errors (incorrectly classified samples) using its predicted uncertainty. Specifically, we define the predicted uncertainty as $1-p_{\text{max}}$, where $p_{\text{max}}$ is the maximum softmax probability assigned to any class. The ROC curve is constructed by plotting the true positive rate against the false positive rate at varying uncertainty thresholds, where incorrect predictions are treated as the positive class (label 1) and correct predictions as the negative class (label 0). The AUROC corresponds to the area under this curve. A higher AUROC indicates that the model assigns larger uncertainty values to its errors than to its correct predictions, reflecting stronger error-detection capability.
(2) Pearson Correlation Between Ground-Truth Measurement Uncertainty and Predicted Uncertainty. A central contribution of this work is introducing a new measure of uncertainty awareness, made possible by the ground-truth noise statistics provided in ModelNet40-E. Each point cloud sample comes with a measurement uncertainty $\sigma$ that reflects the standard deviation of the LiDAR-like noise applied during data generation. This allows us to directly test whether a model’s predicted uncertainty increases when the true measurement uncertainty increases.
To quantify this relationship, we compute the Pearson correlation coefficient between ground-truth measurement uncertainty $X$ and predicted uncertainty $Y$:
$r=\frac{\mathrm{cov}(X,Y)}{\sigma_{X}\,\sigma_{Y}}.$
Predicted uncertainty is defined as $1-p_{\text{max}}$, where $p_{\text{max}}$ is the maximum softmax probability assigned to any class. A high correlation indicates that the model assigns higher uncertainty to noisier inputs, demonstrating strong uncertainty awareness.
However, measuring correlation on all samples mixes two effects: genuine uncertainty awareness and the natural confidence drop under noise. To isolate the true awareness signal, we compute the Pearson correlation using only correctly classified samples, capturing whether the model’s confidence decreases appropriately as noise increases while still predicting the correct class. For completeness, we also report the correlation over all samples for comparison.
4.3 Main Results
4.3.1 Classification Accuracy
Table 2 reports classification accuracy across all models and noise levels. As expected, performance degrades as noise severity increases. Early architectures such as PointNet, PointNet++ and DGCNN experience sharp degradation under noise, whereas more recent methods such as SimpleView and PTv3 maintain a higher accuracy under noise, indicating stronger robustness to perturbations. Notably, SimpleView achieves the lowest accuracy on the clean dataset but the highest accuracy under light, moderate, and heavy noise, highlighting its resilience to corruption.
Table 2: Classification Accuracy (%) for all models across different noise levels. Highest accuracy under each noise severity is highlighted in bold.